Write an equation of the line that passes through the point (2,3) and is perpendicular to the line y=1/2x-1. If possible I need the work too.
Answer by drk(1908)

First you have an equation solved for Y. Look at that and find the slope; m = 1/2.
Second, the key term is "perpendicular". If you see this term, take your slope and apply the negative reciprocal. So our "new", perpendicular slope is m = -2.
Third, you are given a coordinate. Remember that all coordinates are in (X,Y) form. We go to the formula
Y = mX + b and apply the numbers we know. We find the b value.
3 = (-2)(2) + b
3 = -4 + b
7 = b.
Now we have enough for our equation: Y = -2x + 7.